# Notes on Quasinormal Modes of charged de Sitter Blackholes from Quiver Gauge Theories

### v1

@@ 3 QNMs in RN-dS blackhole
where the coefficients $a_{ij}$ are determined from (16) as explained in (34) and we note that as e.g. $a_{01}=0$, some of the expansion coefficients do not appear above. To work out the dS modes and PS modes for $\Lambda M^{2}=0.08,0.06$, we further needed to include a higher-order term $a_{40}t^{4}/q^{4}$, and at $\Lambda M^{2}=0.08$ we had to add terms up to $a_{50}t^{5}/q^{5}$ to reproduce the numerical result of the dS mode $\omega\simeq-0.16i$.
@@ 4 Analytic results
One can check via numerical computations that in Dictionary 1, the PS modes correspond to the poles of the following $\Gamma$-function,
@@
In Dictionary 2, the instanton expansion parameter $t$ becomes small in the near-extremal region. From the numerical computation, we see that the dS modes seem to come from the poles of the following $\Gamma$-function,
@@ 5 Application to non-scalar Perturbations in RN-dS blackhole
where again $n=0,1,2,\dots\in\mathbb{N}$. We note that the leading-order analytic expression again gives better results than the brute-force connection formula computation, which suggests that re-expansion is necessary for the connection formula to write down a series expansion with a clear convergence radius.
@@ 6 Conclusion and Discussions
In this article, we generalized the prescription proposed in Bonelli:2022ten to write down the connection formulae for ordinary differential equations with five or six regular singularities. We checked the effectiveness of the connection formulae by computing the QNMs in RN-dS blackholes in both numerical and analytic (series expansion) ways. The analytic expressions of different families of QNMs can be respectively obtained from different coordinate transformations and dictionaries, based on expansion over different parameters. The leading order behaviors of the dS modes and near-extremal modes reproduce the known result conjectured in the literature, and we also provide the analytic expression for PS modes. Our method can systematically be extended to arbitrary high orders. In general, we also saw that even the leading order analytic form of QNMs is supreme to the brute-force computation of connection formula. This suggests that we need to re-organize or re-expand the connection formula, which was originally based on the instanton expansion in quiver gauge theories, in blackhole parameters. In the Heun equation case, similar techniques have been applied in Lisovyy:2022flm . We wish to provide such a perturbative connection formula and apply it to study the strong cosmic censorship in a purely analytic way in our future work. This will also help to analyze the connection formula in the colliding limit (c.f. a program initiated in Zhao:2024rrw ), where irregular singularities can further be put into the differential equation by colliding several regular singularities at a coincident point.
We also remark that one can go beyond to consider the dS-Kerr-Newman blackhole with both charge and rotation parameters. The radial part of the Teukolsky equation can also be mapped to an ODE with five singularities Suzuki:1998vy (see Appendix C for the dictionary). Due to the complication of the angular and radial equations coupled together, we will leave it to future work to investigate the numerical and analytic results of the QNMs of such complicated blackhole geometries.
@@ Appendix B More connection formulae
Appendix C dS-Kerr-Newman blackhole and 6-pt BPZ equation

### v3

@@ 3 QNMs in RN-dS blackhole
where the coefficients $a_{ij}$ are determined from (16) as explained in (34) and we note that as e.g. $a_{01}=0$, some of the expansion coefficients do not appear above. We remark that many coefficients $a_{ij}$ with $i<j$ are vanishing. This is due to the fact that $a$ is computed from the first relation in (16), and because of the asymmetry between $t/q$ and $q$ introduced by the $t$-derivative $t\partial_{t}$, one can directly see that $a_{0i}$ vanishes for all $i$ due to the $t$-derivative. Interestingly $a_{12}$ further vanishes in a very non-trivial way, but we are not sure whether there is any profound physical reason behind it. To work out the dS modes and PS modes for $\Lambda M^{2}=0.08,0.06$, we further needed to include a higher-order term $a_{40}t^{4}/q^{4}$, and at $\Lambda M^{2}=0.08$ we had to add terms up to $a_{50}t^{5}/q^{5}$ to reproduce the numerical result of the dS mode $\omega\simeq-0.16i$.
@@ 4 Analytic results
One can check via numerical computations that in Dictionary 1, the PS modes correspond to the poles of the following $\Gamma$-function in the factor $\mathcal{M}$ in (29),
@@
We further remark that the analytic expansion in terms of $(1-9\Lambda)^{\frac{1}{2}}$ and $Q^{2}(1-9\Lambda)^{-\frac{1}{2}}$ is a re-expansion of the instanton series in the region $|t/q|\ll|q|<1$. This new series expansion works because the instanton-expansion coefficients before terms involving only $q$, i.e. $a_{0i}$, are all vanishing as discussed before.
@@
In Dictionary 2, the instanton expansion parameter $t$ becomes small in the near-extremal region. From the numerical computation, we see that the dS modes seem to come from the poles of the following $\Gamma$-function appearing in the prefactor $\mathcal{M}$ in the connection formula (133),
@@ 5 Application to non-scalar Perturbations in RN-dS blackhole
where again $n=0,1,2,\dots\in\mathbb{N}$. We note that the leading-order analytic expression again gives even better results than the brute-force connection formula computation including higher-order instanton corrections when $\Lambda M^{2}\approx 0.1$ (see Table 6), which suggests that re-expansion is necessary for the connection formula to write down a series expansion with a clear convergence radius.
Conformal block data QNM package Leading analytic results
$\Lambda M^{2}=0.105$$\pm 0.0815791-0.0235535i$$\pm 0.0614357-0.0231477i$$\pm 0.0771136-0.0232506i$
$s=1$,$l=1$$\pm 0.080743-0.0706604i$$\pm 0.0609648-0.0694504i$$\pm 0.0771136-0.0697518i$
±$\pm 0.0790442-0.117767i$$\pm 0.060001-0.115778i$$\pm 0.0771136-0.116253i$
$\Lambda M^{2}=0.105$$\pm 0.12771-0.0235535i$$\pm 0.1114268-0.0231930i$$\pm 0.120814-0.0232506i$
$s=1$,$l=2$$\pm 0.127178-0.0706604i$$\pm 0.1111326-0.0695827i$$\pm 0.120814-0.0697518i$
±$\pm 0.126106-0.117767i$$\pm 0.11054-0.11598i$$\pm 0.120814-0.116253i$
Table 6: Comparison of QNMs in the RN-dS blackhole obtained from the 6pt connection formula with the leading-order analytic results. We see that although only the leading contribution is taken in the analytic expression, it still gives slightly better numerical results than the brute-force instanton computation.
@@ 6 Conclusion and Discussions
In this article, we generalized the prescription proposed in Bonelli:2022ten to write down the connection formulae for ordinary differential equations with five or six regular singularities. We checked the effectiveness of the connection formulae by computing the QNMs in RN-dS blackholes in both numerical and analytic (series expansion) ways. The analytic expressions of different families of QNMs can be respectively obtained from different coordinate transformations and dictionaries, based on expansion over different parameters. The leading order behaviors of the dS modes and near-extremal modes reproduce the known result conjectured in the literature, and we also provide the analytic expression for PS modes. Our method can systematically be extended to arbitrary high orders. In general, we also saw that even the leading order analytic form of QNMs is supreme to the brute-force computation of connection formula. This suggests that we need to re-organize or re-expand the connection formula, which was originally based on the instanton expansion in quiver gauge theories, in blackhole parameters. In the Heun equation case, similar techniques have been applied in Lisovyy:2022flm . We wish to provide such a perturbative connection formula and apply it to study the strong cosmic censorship in a purely analytic way in our future work (including the instability of the fundamental QNM in dS blackholes in various dimensions Rahman:2018oso ; Rahman:2020guv ; Sarkar:2023rhp ). It will also help to analyze the connection formula in the colliding limit (c.f. a program initiated in Zhao:2024rrw ), where irregular singularities can further be put into the differential equation by colliding several regular singularities at a coincident point. We also remark that one can go beyond to consider the dS-Kerr-Newman blackhole with both charge and rotation parameters. The radial part of the Teukolsky equation can also be mapped to an ODE with five singularities Suzuki:1998vy (see Appendix C for the dictionary). Due to the complication of the angular and radial equations coupled together, we will leave it to future work to investigate the numerical and analytic results of the QNMs of such complicated blackhole geometries.
Acknowledgment
We thank Hongfei Shu for helpful discussions. R.Z. is supported by National Natural Science Foundation of China No. 12105198 and the High-level personnel project of Jiangsu Province (JSSCBS20210709).
@@ Appendix B More connection formulae
Appendix C dS-Kerr-Newman blackhole and 5-pt BPZ equation